# Large Language Model

### v1

@@ Abstract
The event-based Vision-Language Model (VLM) recently has made good progress for practical vision tasks. However, most of these works just utilize CLIP for focusing on traditional perception tasks, which obstruct model understanding explicitly the sufficient semantics and context from event streams. To address the deficiency, we propose EventVL, the first generative event-based MLLM (Multimodal Large Language Model) framework for explicit semantic understanding. Specifically, to bridge the data gap for connecting different modalities semantics, we first annotate a large event-image/video-text dataset, containing almost 1.4 million high-quality pairs of data, which enables effective learning across various scenes, e.g., drive scene or human motion. After that, we design Event Spatiotemporal Representation to fully explore the comprehensive information by diversely aggregating and segmenting the event stream. To further promote a compact semantic space, Dynamic Semantic Alignment is introduced to improve and complete sparse semantic spaces of events. Extensive experiments show that our EventVL can significantly surpass existing MLLM baselines in event captioning and scene description generation tasks. We hope our research could contribute to the development of the event vision community.
1 Introduction
Event cameras are sensors that asynchronously measure changes in intensity at each pixel with microsecond-level precision. Compared to traditional cameras, event cameras offer several significant advantages [14], including a high dynamic range ($>120$ dB), resistance to motion blur, high temporal resolution, and extremely low power consumption.These advantages have led to significant successes of event-based methods in different domains, such as robotics [18] and imaging applications [44, 36]. Even though event cameras offer richer semantic information—including continuous temporal motion and higher dynamic range lighting, exploration into high-level and open-world understanding tasks remains limited.
@@ 1 Introduction
To solve these challenges mentioned above, we propose EventVL, the first event-based MLLM framework, as shown in Figure 2. As for data engineering, we first annotate almost 1.4 million high-quality image-text-event paired data by utilizing current powerful open-source MLLM models [7], which has surpassed the commercial MLLMs like GPT-4o  [1] in many understanding benchmarks. As for EventVL, we propose Event Spatiotemporal Representation to further excavate the spatiotemporal correlation within event data. We spatially and temporally segment event data into different frames and leverage attention blocks to capture the semantic interaction, which can enhance feature expression. To fine-grain align the latent feature space for images and events, we propose the Dynamic Semantic Alignment for rigorous high-resolution semantics extraction, enabling fine-grained alignment between events and images.
Figure 1: The Designed Description Generation Engine. For different kinds of datasets, we pre-define the corresponding problem list and prompts, which “Static Images” is for N-ImageNet [23], N-Caltech101 [31], “Human Motions” is for HARDVS [40] and “Drive Scenes” is for DSEC [16]. The coarse captions generated by MLLMs also are checked manually by random sampling. The pairs data marked as bad descriptions will be used to regenerate descriptions again with modified problems.
@@
Comprehensive experiments verify that our proposed EventVL surpasses other SOTAs in zero/few-shot event captioning, event stream description generation tasks, and event-based dialogue. Furthermore, with a tiny number of parameters (almost 2.3B) compared to other MLLMs [30, 38], our model enables a low-cost for deployment to the real world. As shown in Figure LABEL:fig:abstract (a), EventVL can fully understand semantics and generate precise descriptions for event streams across diverse domains. Additionally, it supports multi-turn dialogues based on event data, enabling deeper semantic understanding, as shown in Figure LABEL:fig:abstract (b). In summary, our contribution has three folds:
@@
(1) We present the First Event-based MLLM Framework, named EventVL, which aligns the large-scale event encoder with the image encoder and LLM. The model demonstrates strong performance on event-based generative tasks such as description or dialogue.
(2) We propose a simple but efficient description generation engine to get high-quality multi-modalities data by utilizing open-source VLMs. The processed datasets contain almost 1.4 million paired data across various domains. As far as we know, it is the biggest multi-modalities pair dataset in the event-based community.
@@
(3) Event Spatiotemporal Representation is proposed for feature adaptive and efficient aggregation. We also propose a Dynamic Semantic Alignment module for fine-grained feature alignment, resulting in a precise and comprehensive event-based understanding.
2 Related Work
@@
Event-based VLMs: Unlike the well-established field of image understanding, the event community is in the early stages of developing universal models for event comprehension. Initial efforts [24, 47] focused on using self-supervision to create a pre-trained backbone for event understanding. Inspired by CLIP’s success in computer vision, researchers began to explore its application for detailed and fine-grained event analysis. [43] suggested transforming event streams into an RGB-style representation to achieve event-text alignment using CLIP’s architecture. Recently, CLIP-based approaches have been applied to more complex downstream tasks such as motion classification [40] and scene segmentation [25]. However, these studies have mainly concentrated on discriminative tasks, falling short of achieving fine-grained understanding, which restricts their application in generative tasks like interactive event-conditioned text generation. In response to the increasing need for a nuanced understanding of event streams, we are committed to introducing a unified event-based MLLM framework. This framework is designed to fully comprehend the semantics of event streams through explicit language and is compatible with various types of event data.
Open-Source MLLMs: The emergence of open-source MLLMs [29, 17, 38, 30] has profoundly impacted the field of Artificial General Intelligence (AGI) by bolstering the ability to process both visual and textual data. In the past year, several open-source MLLMs have gained prominence such as MiniGPT-4 [53] and VisionLLM [39]. However, these models typically rely on traditional images or videos as input, which can result in the generation of low-quality images under extreme imaging conditions, thereby impairing comprehensive world understanding. Event cameras, with their distinct properties, can effectively address these challenges. Therefore, we are exploring the integration of event-based data into existing MLLMs to unlock the full semantic potential of event data.
3 Data Engineering
@@
Figure 2: Overview of our proposed EventVL. EventVL mainly consists of two parts: (a) Event Spatiotemporal Representation and (b) Dynamic Semantic Alignment. First, we introduce Event Spatiotemporal Representation for extracting diverse event information comprehensively, which consists of two split techniques –(i) Hierarchical Temporal Split and (ii) Adaptive Spatial Split. Then, we utilize Dynamic Semantic Alignment to further push the image and event feature space in fine-grained alignment.
A high-quality dataset is crucial for training a MLLM. Some datasets pairing images or videos with events [23, 40] offer only basic categorical text annotations, akin to ”This is a category”, which lack detail and variety. While certain studies suggest using coarse image-text data to address the limited diversity of image-event datasets like N-ImageNet, these annotations often have a weak connection to the images and include much extraneous information. Our experiments indicate that training with such datasets leads to a decline in performance. Therefore, it is imperative to re-annotate this data to create a more precise and granular dataset, thereby enhancing the MLLM’s comprehension.
As shown in Figure 1, we design a simple but effective generation engine for different domain datasets that contain static images, dynamic scenes and human motions. We chose the InternVL2-76B [7] as the foundation model for description generation, which has shown its superiority and surpasses several SOTAs such as GPT-4o or Claude 3.5 on various understanding benchmarks. Utilizing it to label our data can reduce multiple costs compared to commercial engines and bring higher-quality descriptions.
Coarse Generation. We manually design the prompts and problem lists corresponding to various datasets, which can be referred to in our appendix. For image-event datasets (N-ImageNet, N-Caltech101), we first randomly sample questions from the problem list, then directly send images with questions into the InternVL2-76B for generating coarse captions. For video-event datasets (HARDVS, DSEC), it is inefficiency that utilizes all frames for generating description, which causes a large amount of computational consumption [25, 8]. To effectively extract long-term information from origin videos, we sampled uniformly $n$ frames. If the video contains less than $n$ frames, we adopt all frames. For balancing the cost and performance, we chose $n=14$. Then, the video and the correspondiong instruct is sent to InternVL2-76B for getting detailed description.
Manual Check. The cost of manually checking every image/video-text description is enormous. Hence, we develop the sampling strategies based on the category classification of each dataset for checking. For N-ImageNet, N-Caltech101, HARDVS and DSEC, they contain 1000 classes, 101 classes, 300 classes and 20 scenes respectively. We sample five coarse descriptions from each class or scene for checking. If the output description is not so good, all images/videos from this corresponding class or scenes will be sent to MLLM to regenerate the captions again with modified prompts and designated problems. Finally, all annotated multimodal pair datasets except N-Caltech101 will be mixed for model training. The dataset performance is validated as shown in Table 4. Note that our prepared dataset has almost 1.4 million image/video-event-text pairs covering diverse domains, which guarantees the event understanding performance.
4 Method
@@
4.1 Event Spatiotemporal Representation
@@
Adaptive Spatial Split. Previous research has focused on encoding complete event representations and aligning them with corresponding images within the feature space. While this method has led to improved model performance [48, 46], it requires resizing event frames to a low resolution, such as $224\times 224$. This resizing results in a significant loss of spatial information, particularly for high-resolution events like those in N-ImageNet [23], which can measure $480\times 640$. This issue worsens under poor imaging conditions, such as overexposure or rolling shutter effects, leading to even greater information loss. Hence, Adaptive Spatial Split is introduced for exploring full spatial event-based category information, as shown in Figure 2 (ii), which adaptively splits event representation or images into multiple patches uniformly in a high-resolution manner.
To preserve natural aspect ratios during processing, we adaptively select the most suitable aspect ratio from a predefined set $\mathcal{K}$, which encompasses all possible aspect ratio combinations derived from $n_{min}$ to $n_{max}$ tiles. We set $n_{min}$ to 1 and set $n_{max}$ to 6. For each input image or event representation, we determine its aspect ratio and compare it with the predefined ratios by calculating the absolute difference. In cases where multiple predefined aspect ratios are close matches (e.g., 1:1 and 2:2), we prioritize the one that does not exceed twice the area of the event representation to avoid significant enlargement. Subsequently, we divide the event representation into patches according to this chosen ratio and resize these patches to the desired dimensions. In detailed, given the multi-level events $\hat{\bm{x}}_{ev}=\{\{\hat{x}^{r}_{ev1}\}^{N_{1}}_{r=1},\{\hat{x}^{r}_{ev2}\}^{N_{2}}_{r=1},\hat{x}_{ev3}\}$, we only apply Adaptive Spatial Split to the level-3 event frame $\hat{x}_{ev3}$ and obtain multiple event patches spatially $\{\{\hat{x}^{r}_{ev3}\}^{N_{p}}_{r=1}\}$ ($N_{p}=6$). We concatenate multiple event patches with the origin multi-level events, obtaining Event Spatiotemporal Representation $\tilde{\bm{x}}_{ev}=\{\{\hat{x}^{r}_{ev1}\}^{N_{1}}_{r=1},\{\hat{x}^{r}_{ev2}\}^{N_{2}}_{r=1},\hat{x}_{ev3},\{\hat{x}^{r}_{ev3}\}^{N_{p}}_{r=1}\}$.
4.2 Dynamic Semantic Alignment
@@
$\vspace{-0.1cm}\mathcal{L}_{c}=1-\mathop{\textnormal{cos}}(\bm{\phi}_{ev},\bm{\phi}_{im})=1-\frac{\bm{\phi}_{ev}\cdot\bm{\phi}_{im}}{\|\bm{\phi}_{ev}\|\|\bm{\phi}_{im}\|}.$ (1)
@@
Table 1: Statistical information about datasets annotated by Section 3. Each data pair contains an accurate description.
4.3 Framework Optimization
$\vspace{-0.1cm}p(\bm{x}_{t}\mid\bm{\phi}_{ev},\bm{x}_{\textnormal{instruct}})=\prod_{l=1}^{L}\phi_{t}(x_{t}^{l}\mid\bm{\phi}_{ev},\bm{x}_{\textnormal{instruct}},x_{t}^{1:l-1}),$ (2)
$\vspace{-0.1cm}\mathcal{L}_{ev,t}=-\mathop{\textnormal{log}}p(\bm{x}_{t}\mid\bm{\phi}_{ev},\bm{x}_{\textnormal{instruct}}).$ (3)
$\vspace{-0.1cm}\mathcal{L}_{ev,im,t}=-\mathop{\textnormal{log}}p(\bm{x}_{t}\mid\frac{1}{2}(\bm{\phi}_{ev}+\bm{\phi}_{im}),\bm{x}_{\textnormal{instruct}}).\vspace{-0.1cm}$ (4)
The overall training objectives can be written as follows:
$\vspace{-0.1cm}\mathcal{L}=\frac{1}{2}\lambda_{1}(\mathcal{L}_{ev,t}+\mathcal{L}_{ev,im,t})+\lambda_{2}\mathcal{L}_{c}.\vspace{-0.1cm}$ (5)
@@ 4.3 Framework Optimization
N-Caltech101 [31] Qwen2-VL-Chat [38] 10B 0.346 0.153 0.069 0.033 0.096 0.287 0.092
@@
EventVL-Base (Ours) 2.3B 0.360 0.238 0.163 0.115 0.145 0.377 0.823
@@
EventVL-Base (Ours) 2.3B 0.605 0.513 0.450 0.396 0.314 0.628 2.603
N-ImageNet [23] Qwen2-VL-Chat [38] 10B 0.197 0.087 0.040 0.021 0.067 0.208 0.152
@@
EventVL-Base (Ours) 2.3B 0.561 0.392 0.279 0.205 0.167 0.462 0.573
Table 2: Quantitative performance of zero-shot event-based caption generation on different datasets. In this experiment, we only consider the event data as input for accurately evaluating the model’s event-based understanding ability. Our EventVL achieves competitive prediction performance compared with state-of-the-art MLLMs. Gray color denotes the video-based MLLMs.
Val set $n$ Bleu-4 METEOR ROUGE-L CIDEr
N-ImageNet 1 0.056 0.098 0.299 0.387
5 0.078 0.118 0.329 0.550
10 0.089 0.127 0.345 0.646
15 0.090 0.127 0.348 0.683
20 0.095 0.132 0.353 0.720
HARDVS 1 0.212 0.203 0.519 1.205
5 0.328 0.268 0.581 1.970
10 0.345 0.291 0.598 2.129
15 0.369 0.306 0.615 2.364
20 0.392 0.311 0.628 2.604
Table 3: Quantitative results about few-shot caption generation. “$n$” denotes the sample number per category used in training.
4.4 Inference
EventVL offers two kinds of inference pipelines. The first one is to only utilize event modality as input. The second type is to incorporate the image and event modality as input, in which we directly add event embeddings with image embeddings as the input embedding.
Figure 3: Qualitative results about zero-shot caption generation by utilizing N-ImageNet [23] and HARDVS [40] val set. Red color denotes model wrong description and orange color denotes the accurate object description. Underline text means redundant description.
5 Experiments
5.1 Experiments Setup
Implementation details can be referred to appendix.

### v2

@@ Abstract
The event-based Vision-Language Model (VLM) recently has made good progress for practical vision tasks. However, most of these works just utilize CLIP for focusing on traditional perception tasks, which obstruct model understanding explicitly the sufficient semantics and context from event streams. To address the deficiency, we propose EventVL, the first generative event-based MLLM (Multimodal Large Language Model) framework for explicit semantic understanding. Specifically, to bridge the data gap for connecting different modalities semantics, we first annotate a large event-image/video-text dataset, containing almost 1.4 million high-quality pairs of data, which enables effective learning across various scenes, e.g., drive scene or human motion. After that, we design Event Spatiotemporal Representation to fully explore the comprehensive information by diversely aggregating and segmenting the event stream. To further promote a compact semantic space, Dynamic Semantic Alignment is introduced to improve and complete sparse semantic spaces of events. Extensive experiments show that our EventVL can significantly surpass existing MLLM baselines in event captioning and question-answering tasks. We hope our research could contribute to the development of the event community.
I Introduction
Event cameras are sensors that asynchronously measure changes in intensity at each pixel with microsecond-level precision. Compared to traditional cameras, event cameras offer several significant advantages [1], including a high dynamic range (>120>120 dB), resistance to motion blur, high temporal resolution, and extremely low power consumption.These advantages have led to significant successes of event-based methods in different domains, such as robotics [2] and imaging applications [3, 4]. Even though event cameras offer richer semantic information—including continuous temporal motion and higher dynamic range lighting, exploration into high-level and open-world understanding tasks remains limited.
@@ I Introduction
To solve these challenges mentioned above, we propose EventVL, the first event-based MLLM framework, as shown in Figure 1. As for EventVL, we propose Event Spatiotemporal Representation to further excavate the spatiotemporal correlation within event data. We spatially and temporally segment event data into different frames and leverage attention blocks to capture the semantic interaction, which can enhance feature expression. To fine-grain align the latent feature space for images and events, we propose the Dynamic Semantic Alignment for rigorous high-resolution semantics extraction, enabling fine-grained alignment between events and images. As for data engineering, we first annotate almost 1.4 million high-quality image-text-event paired data by utilizing current powerful open-source MLLM models [17], which has surpassed the commercial MLLMs like GPT-4o  [14] in many understanding benchmarks.
@@
Comprehensive experiments verify that our proposed EventVL surpasses other SOTAs in zero/few-shot event captioning, event stream description generation tasks, and event-based dialogue. Furthermore, with a tiny number of parameters (almost 2.3B) compared to other MLLMs [18, 19], our model enables a low-cost for deployment to the real world. In summary, our contribution has three folds:
@@
  • We present the first Event-based MLLM framework, named EventVL, which aligns the large-scale event encoder with the image encoder and LLM. The model demonstrates strong performance on event-based generative tasks such as description or dialogue.
@@
  • Event Spatiotemporal Representation is proposed for feature adaptive and efficient aggregation. We also propose a Dynamic Semantic Alignment module for fine-grained feature alignment, resulting in a precise and comprehensive event-based understanding.
  • We propose the high quality event-image-text paired datasets, which contain almost 1.4 million paired data across various domains. As we known, it is the biggest multi-modalities pair dataset in the event community.
II Related Work
@@
Event-based VLMs. Unlike the well-established field of image understanding, the event community is in the early stages of developing universal models for event comprehension. Initial efforts [20, 21] focused on using self-supervision to create a pre-trained backbone for event understanding. Inspired by CLIP’s success in computer vision, researchers began to explore its application for detailed and fine-grained event analysis. [12] suggested transforming event streams into an RGB-style representation to achieve event-text alignment using CLIP’s architecture. Recently, CLIP-based approaches have been applied to more complex downstream tasks such as motion classification [16] and scene segmentation [9]. However, these studies have mainly concentrated on discriminative tasks, falling short of achieving fine-grained understanding, which restricts their application in generative tasks like interactive event-conditioned text generation. In response to the increasing need for a nuanced understanding of event streams, we are committed to introducing a unified event-based MLLM framework. This framework is designed to fully comprehend the semantics of event streams through explicit language and is compatible with various types of event data.
@@
Figure 1: Overview of our proposed EventVL. EventVL mainly consists of two parts: (a) Event Spatiotemporal Representation and (b) Dynamic Semantic Alignment. First, we introduce Event Spatiotemporal Representation for extracting diverse event information comprehensively, which consists of two split techniques –(i) Hierarchical Temporal Split and (ii) Adaptive Spatial Split. Then, we utilize Dynamic Semantic Alignment to further push the image and event feature space in fine-grained alignment. Finally, we perform Progressive Training strategy to make EventVL better.
III Method
@@
Algorithm 1 Generate Adaptive Ratios
Input: $n_{\text{min}},n_{\text{max}}$ are the min./max. number of patch.
Output: $\mathcal{K}$ is the adaptive ratio set.
[…]
2:for $n\leftarrow n_{\text{min}}$ to $n_{\text{max}}$ do
3:for $i\leftarrow 1$ to $n$ do
4:   for $j\leftarrow 1$ to $n$ do
5:    if $i\times j\leq n_{\text{max}}$ and $i\times j\geq n_{\text{min}}$ do
6:     $\mathcal{K}\leftarrow\mathcal{K}\cup\{(i,j)\}$
[…]
8:return $\mathcal{K}$
III-A Event Spatiotemporal Representation
@@
Adaptive Spatial Split. Previous research has focused on encoding complete event representations and aligning them with corresponding images within the feature space. While this method has led to better performance [8, 22], it requires resizing event frames to a low resolution, such as $224\times 224$. This resizing results in a significant loss of spatial information, particularly for high-resolution events like those in N-ImageNet [15], which can measure $480\times 640$. This issue worsens under poor imaging conditions, such as overexposure or rolling shutter effects, leading to even greater information loss. Hence, Adaptive Spatial Split is introduced to explore full spatial event-based category information, as shown in Figure 1 (ii), which adaptively splits event representation or images into multiple patches uniformly in a high-resolution manner.
To preserve natural aspect ratios during processing, we adaptively select the most suitable aspect ratio from a predefined set $\mathcal{K}$, which encompasses all possible aspect ratio combinations derived from $n_{\min}$ to $n_{\max}$ tiles. We set $n_{\min}$ to 1 and set $n_{\max}$ to 6 as shown in Algorithm 1 to complete the set. For each input image or event representation, we determine its aspect ratio and compare it with the predefined ratios by calculating the absolute difference. In cases where multiple predefined aspect ratios are close matches (e.g., 1:1 and 2:2), we prioritize the one that does not exceed twice the area of the event representation to avoid significant enlargement. Subsequently, we divide the event representation into patches according to this chosen ratio and resize these patches to the desired dimensions. In detailed, given the multi-level events $\hat{\bm{x}}_{ev}=\{\{\hat{x}^{r}_{ev1}\}^{N_{1}}_{r=1},\{\hat{x}^{r}_{ev2}\}^{N_{2}}_{r=1},\hat{x}_{ev3}\}$, we only apply Adaptive Spatial Split to the level-3 event frame $\hat{x}_{ev3}$ and obtain multiple event patches spatially $\{\{\hat{x}^{r}_{ev3}\}^{N_{p}}_{r=1}\}$ ($N_{p}=6$). We concatenate multiple event patches with the origin multi-level events, obtaining Event Spatiotemporal Representation $\tilde{\bm{x}}_{ev}=\{\{\hat{x}^{r}_{ev1}\}^{N_{1}}_{r=1},\{\hat{x}^{r}_{ev2}\}^{N_{2}}_{r=1},\hat{x}_{ev3},\{\hat{x}^{r}_{ev3}\}^{N_{p}}_{r=1}\}$.
III-B Dynamic Semantic Alignment
@@
$\mathcal{L}_{c}=1-\mathop{\textnormal{cos}}(\bm{\phi}_{ev},\bm{\phi}_{im})=1-\frac{\bm{\phi}_{ev}\cdot\bm{\phi}_{im}}{\|\bm{\phi}_{ev}\|\|\bm{\phi}_{im}\|}.$ (1)
@@
III-C Framework Optimization
$\vskip-2.84544ptp(\bm{x}_{t}\mid\bm{\phi}_{ev},\bm{x}_{\textnormal{instruct}})=\prod_{l=1}^{L}\phi_{t}(x_{t}^{l}\mid\bm{\phi}_{ev},\bm{x}_{\textnormal{instruct}},x_{t}^{1:l-1}),$ (2)
$\mathcal{L}_{ev,t}=-\mathop{\textnormal{log}}p(\bm{x}_{t}\mid\bm{\phi}_{ev},\bm{x}_{\textnormal{instruct}}).$ (3)
$\mathcal{L}_{ev,im,t}=-\mathop{\textnormal{log}}p(\bm{x}_{t}\mid\frac{1}{2}(\bm{\phi}_{ev}+\bm{\phi}_{im}),\bm{x}_{\textnormal{instruct}}).$ (4)
The overall training objectives can be written as follows:
$\mathcal{L}=\frac{1}{2}\lambda_{1}(\mathcal{L}_{ev,t}+\mathcal{L}_{ev,im,t})+\lambda_{2}\mathcal{L}_{c}.$ (5)
III-D Progressive Training and Inference
We adopt a two-stage approach for model training. In the first stage, we fine-tune the event encoder while keeping all other modules frozen, utilizing the EventVL-Base dataset. This step is designed to enhance the foundational event comprehension ability of the image-based encoder. In the second stage, we fine-tune the text decoder while freezing the other modules, using the EventVL-QA dataset to improve the model’s fine-grained analytical capabilities for event-text pairs. This progressive training strategy enables us to develop a world model that comprehensively understands events across various scenarios. Furthermore, EventVL supports two inference pipelines. The first pipeline processes only the event modality as input, while the second integrates both image and event modalities. In the latter, event embeddings are directly combined with image embeddings to form the input representation, obtaining more comprehensive understanding of the event.
TABLE I: Statistical information about EventVL dataset annotated by Section IV. Each data pair contains an accurate description.
@@ III-D Progressive Training and Inference
TABLE II: Statistical information about our proposed EventVL-QA dataset. Each data pair contains an accurate answer.
Dataset Type Resolution Split Category Scale
N-ImageNet [15] Image $480\times 640$ train 160 16,000
val 40 4,000
HARDVS [16] Video $346\times 260$ train 480 9,600
val 120 2,400
DSEC [25] Video $640\times 480$ train 42 43,460
Total - - - 842 75,460
IV Data Engineering
A high-quality dataset is crucial for training a MLLM. Some datasets pairing images or videos with events [15, 16] offer only basic categorical text annotations, akin to ”This is a category”, which lack detail and variety. While certain studies suggest using coarse image-text data to address the limited diversity of image-event datasets like N-ImageNet, these annotations often have a weak connection to the images and include much extraneous information. Hence, we adopt two-stage pipeline to generate the high-quality annotation for the selected datasets: N-ImageNet [15], N-Caltech [26], HARDVS [16] and DSEC [25]. In detailed, we first utilize the powerful MLLM engine like InternVL2-76B to generate the coarse caption, and then check it manually in random sampling manner. Finally, we construct the EventVL-Base dataset as shown in Table I, comprising 1,393,444 meticulously annotated multimodal pairs with rich event descriptions. Additionally, we introduce EventVL-QA as shown in Table II, a high-quality dataset containing 75,460 QA pairs designed to enhance model dialogue and comprehension capabilities. Specifically, EventVL-QA includes multiple QA pairs per category. For N-ImageNet, HARDVS, and DSEC, we define 10, 20, and 20 carefully curated sequential questions per category, respectively.
V Experiments
TABLE III: Quantitative performance of zero-shot event-based caption generation on different datasets. In this experiment, we only consider the event data as input for accurately evaluating the model’s event-based understanding ability. Our EventVL achieves competitive prediction performance compared with state-of-the-art MLLMs. Gray color denotes the video-based MLLMs. InternVL-E denotes the InternVL finetuned by our proposed datasets.
@@ V Experiments
N-ImageNet [15] Qwen2-VL-Chat [19] 10B 0.346 0.153 0.069 0.033 0.096 0.287 0.092
@@
InternVL-E [13] 2.3B 0.360 0.233 0.144 0.092 0.138 0.354 0.423
EventVL (Ours) 2.3B 0.367 0.248 0.183 0.125 0.157 0.397 0.853
@@
InternVL-E [17] 2.3B 0.457 0.322 0.289 0.245 0.211 0.498 1.845
EventVL (Ours) 2.3B 0.623 0.534 0.464 0.407 0.326 0.637 2.619
N-Caltech101 [26] Qwen2-VL-Chat [19] 10B 0.197 0.087 0.040 0.021 0.067 0.208 0.152
@@
InternVL-E [13] 2.3B 0.421 0.298 0.163 0.144 0.328 0.345 0.411
EventVL (Ours) 2.3B 0.574 0.408 0.283 0.223 0.183 0.477 0.599
V-A Implementation
General Setting. In Dynamic Semantic Alignment, we set $n_{\min}=1$ and $n_{\max}=6$. Additionally, we set $\lambda_{1}=1$ and $\lambda_{2}=1$ to balance the multimodal alignment. Due to the differences in event point scales between N-ImageNet and N-Caltech101, we set distinct event point numbers per frame, $n_{\epsilon}$. Specifically, $n_{\epsilon}=40000$ for N-ImageNet and $n_{\epsilon}=20000$ for N-Caltech101. If the total event point number, $N_{\epsilon}$, does not meet the requirements for all levels of event-based representation from Section III-A, we pad the event stream during post-processing. Note that we do not use Event Spatiotemporal Representation for video-event pairs from the HARDVS [16] and DSEC [25] datasets, as they contain sufficient temporal and spatial information. Instead, we directly use cosine similarity from Equation 1 to align the video-event modalities. The final optimization objectives remain the same as for the image-event pair datasets. Due to the imbalanced data scales across domains, we adopt a weighted sampling strategy for hybrid training. We set the sampling weights as follows: 0.6 for N-ImageNet, 0.1 for DSEC, and 0.3 for HARDVS. All training is conducted on 8 Nvidia GTX A6000 GPUs for approximately 144 hours. We set the batch size to 2 and use the DeepSpeed  [^1] https://github.com/microsoft/DeepSpeed ZeRO-1 training strategy to optimize training efficiency. For InternVL-E, we fine-tune InternVL2-2B using all proposed datasets for 1 epoch.
Zero-shot Setting. We use the N-ImageNet, HARDVS, and DSEC train splits for training. To prevent overfitting, we perform training for 1 epoch. We evaluate EventVL on the validation sets for zero-shot description generation, using only event data as input.
Few-shot Setting. For the few-shot experiments, we sample $n$-shot data from the entire N-ImageNet, HARDVS, and DSEC datasets for training. The remaining datasets, including N-Caltech101, are used for evaluation. We perform all few-shot training for 3 epochs to ensure sufficient learning of event-based category knowledge.
V-B Datasets